When 1 mol of sodium nitrate, is dissolved in 1 cubic decimeter of water, 40 kJ of heat energy is absorbed. What is the drop in temperature when 17.0g of sodium nitrate is dissolved in 1 cubic decimeter of water? [Relative atomic mass:N=14, O=16,Na=23, specific heat capacity of water=4J/g°Cand density of solution = 1.0 gcm-3] ]

Answers option:
a) 4°C
b) 2°C
c) 1°C
d) 5°C
1
Expert's answer
2020-05-19T08:47:21-0400

Q=mc∆T.

∆T= Q/mc= 40 x 17/85 x 17 x 4 = 0.1.

c) 1°C